2.6. If G is a group in which (ab)i = aibi for three consecutive integers i for all a, b ∈ G, show that G is abelian.


1
Expert's answer
2022-06-21T06:42:33-0400

Observe that if there exist two consecutive integers "n" , "n+1" such that "(ab)^n=a^nb^n" and "(ab)^{n+1}=a^{n+1}b^{n+1}" for all "a,b\in G", then "a^{n+1}b^{n+1}=(ab)^{n}ab=a^nb^nab". Then we obtain "a^{n+1}b^{n+1}=a^nb^nab". Now by multiplying this equation from left by "a^n" and right by "b^{-1}" we obtain "ab^n=b^na".

In our case taking "n=i" and "n=i+1", we have "ab^i=b^ia" and by taking "n=i+1" and "i+2" we have "ab^{i+1}=b^{i+1}a".

This shows that "ab^{i+1}=b^{i+1}a=bb^ia=bab^i" and now multiplying from right by "b^i" we obtain "ab=ba". Hence "G" is abelian